OpenStudy (anonymous):

The value 5 is a solution for . x^2-6x+5=0 A. True B. False

OpenStudy (lalaly):

true

OpenStudy (anonymous):

A. True

OpenStudy (anonymous):

just substitute and see (5)² -6(5) + 5 = 25 - 30 + 5 = 30 - 30 = 0